OpenStudy (bluemoon1):

Write the slope- intersept form of the equation of each line. 7x+2y=2 x-7y=-42

OpenStudy (bluemoon1):

Those are two separate questions

OpenStudy (bluemoon1):

@amorfide since you're so amazing can you help

OpenStudy (bluemoon1):

Would it be y = -7/2x+1?

OpenStudy (sailorsoju):

2y = -7x+2 y = -(7/2)x + 1

OpenStudy (bluemoon1):

It wouldn't be negative?

OpenStudy (bluemoon1):

Oh never mind

OpenStudy (sailorsoju):

it would be! -(7/2) is negative

OpenStudy (bluemoon1):

Miss read it sorry. How about the second one?

OpenStudy (sailorsoju):

x-7y=-42 -7y = -x -42 y = (1/7)x + 6

OpenStudy (bluemoon1):

Thank you!

OpenStudy (sailorsoju):

i hope that helped, you're welcome :)

OpenStudy (bluemoon1):

Can you help with more?

OpenStudy (bluemoon1):

Write the slope intercept form of the equation of the line through the given point with the given slope. Through (-3,3), slope = 1/3 @sailorsoju